# A Unified, Practical, and Understandable Summary of Non-transactional Consistency Levels in Distributed Replication

### v1

@@ 4.1. Linearizability
Protocols & Systems. Linearizability is the predominant consistency level adopted by critical replication systems built atop SMR protocols. Classic protocols include Chain Replication (Renesse and Schneider, 2004), Multi-Paxos (Lamport, 2001) and its many variants/optimizations (Lamport and Massa, 2004; Lamport, 2006; Lamport et al., 2009; Howard et al., 2016; Oki and Liskov, 1988; Ongaro and Ousterhout, 2014; Mao et al., 2008; Moraru et al., 2013; Poke et al., 2017; Ailijiang et al., 2020; Enes et al., 2020; Bezerra et al., 2014; Stathakopoulou et al., 2022; Whittaker et al., 2021; Ganesan et al., 2020, 2021; Ngo et al., 2020; Mu et al., 2014; Wang et al., 2020; Fouto et al., 2022), Byzantine fault-tolerant protocols (Castro et al., 1999; Yin et al., 2019; Abd-El-Malek et al., 2005; Cowling et al., 2006), and others (Lee et al., 2015; Pan et al., 2021; Li et al., 2016; Takruri et al., 2020; Ports et al., 2015; Zhang et al., 2015) (some with advanced hardware assumptions). Systems incorporating SMR components include lock/coordination services (Burrows, 2006; Hunt et al., 2010; Balakrishnan et al., 2013a, b), replication engines in distributed databases (Corbett et al., 2013; VanBenschoten et al., 2022; Huang et al., 2020; ScyllaDB, 2023; Zhou et al., 2021; Verbitski et al., 2017a; Dragojević et al., 2015), and metadata services of large-scale storage systems (Bolosky et al., 2011; Ghemawat et al., 2003; etcd, 2023; Guerraoui et al., 2022).
4.2. Sequential Consistency
Sequential consistency was first defined by Lamport in (Lamport, 1979). The original interpretation of this level was that all clients agree on the same sequence of operations applied by the service, where operations from each client appear in the sequence in the same order as they are issued by the client, hence the name “sequential consistency”. In our model, a service that provides sequential consistency always gives an ordering that is SO and CASL for any physical timeline workload.
Compared to linearizability, since the ordering does not have to be RT anymore, sequential consistency allows the service to move an operation (or a group of operations) backward in time, reordering it before another group that does not causally precede it. This property is sometimes referred to as unstable ordering (Bessani et al., 2010; Burke et al., 2020), in contrast to stable ordering provided by linearizability.
For example, given the following physical timeline:
$c:$$d:$$c\text{W}x{\angle}1$$d\text{W}x{\angle}2$$d\text{R}x$
A linearizable ordering must be SO and RT:
$|c\text{W}x{\angle}1|$ $|d\text{W}x{\angle}2|$ $|d\text{R}x{:}2|$
$|d\text{W}x{\angle}2|$ $|c\text{W}x{\angle}1|$ $|d\text{R}x{:}1|$
At first glance, it may be hard to tell the exact differences between linearizability and sequential consistency. Attiya and Welch presented a quantitative analysis of the performance implications of these two levels, showing that linearizability is strictly more expensive to implement than sequential consistency for common object types in systems without perfectly synchronized clocks (Attiya and Welch, 1994). But what semantic power do we lose if we relax the real-time guarantee and allow certain reordering? The following paragraphs explain their three practical differences: 1) sequential consistency does not capture external causality dependencies, 2) sequential consistency is non-local, and 3) it takes extra care to add read-modify-write (RMW) operation support to a sequentially-consistent protocol.
External Causality Dependencies. So far we have assumed that all clients communicate only with the service and there are no external communication channels between clients that bypass the service, as depicted in Figure 1. However, in real distributed systems such as cloud databases (Kossmann et al., 2010; Verbitski et al., 2017b; Dageville et al., 2016; Guo et al., 2022), clients of a replicated storage service may be part of a higher-level system. It is common for clients to coordinate with each other through external causality dependencies, which are impossible for the service to capture without using physical time.
In the example depicted by Figure 3, it could be that client $c$ first issues a write of value 1 to object $x$ and waits for its acknowledgment. It then sends a message to client $d$ through an external inter-client channel saying “I have finished my write to $x$ and you can go ahead to operate on $x$”. Client $d$ then issues its own write of value 2 and expects to read out value 2 afterwards. However, since the message from $c$ to $d$ is external to the service, a sequentially consistent service may reorder $d$’s write ahead of $c$’s, and return value 1 for $d$’s read.
@@ 4.2. Sequential Consistency
A service that provides linearizability will be able to capture such implicit external dependencies, because of the real-time property (as $|d\text{W}x{\angle}2|$ starts after $|c\text{W}x{\angle}1|$’s acknowledgment in physical time). We say that linearizability captures external causality dependencies [^5] Note that this is not to be confused with the external consistency property in distributed transaction processing implementations (Gifford, 1981; Corbett et al., 2013), which means that transactions are processed in the same order as they commit., while sequential consistency does not.
Implementation Locality. Herlihy and Wing have proven in (Herlihy and Wing, 1990) that a protocol that implements sequential consistency for each object does not necessarily guarantee overall sequential consistency across all operations. Formally, we say that sequential consistency is non-local: it is possible for an ordering to be SO and CASL on each object, while not SO or CASL overall.
@@
Adding support for read-modify-write (RMW) operations to such protocols is not a trivial task (Burke et al., 2020). In particular, we cannot simply treat RMW operations in the same way as pure writes, because RMWs require a stable base value to determine the result of the read. Systems that demand compare-and-swap (CAS) operations, such as the LogOnce operation in (Guo et al., 2022), may have to go for a service that provides linearizability (or regular sequential consistency (Helt et al., 2021), as will be described in Section 4.5.1).
Protocols & Systems. Despite its significance in memory consistency theory (Lamport, 1979; Adve and Gharachorloo, 1996; Hill, 1998), distributed replication protocols rarely implement the vanilla form of sequential consistency due to aforementioned drawbacks to linearizability and the fact that it offers no significant performance implications. Sequential consistency instead often appears in its transactional form, i.e., serializability (Berenson et al., 1995), or in its per-key form (Cooper et al., 2008) (see Section 4.5.6).
@@ 4.3. Causal+ Consistency
If a global total order is not required, it may be desirable to further relax sequential consistency and embrace the family of causal consistency levels. Causal consistency stems from the definition of causal memory (Ahamad et al., 1995). Lloyd et al. pointed out in (Lloyd et al., 2011) that distributed replication protocols typically implement a slightly stronger version of causal consistency, which they term causal+ consistency. It is essentially causal consistency with convergence guarantee.
In our model, a service that provides causal+ consistency always gives an ordering that is CPO and CASL. Compared to sequential consistency, the ordering does not have to be a serial total order but instead may leave certain operations unordered with each other. This opens opportunities to improve the scalability of a replication protocol. However, all causal dependencies still have to be reflected in the decided ordering.
@@
Notice that $|c\text{W}x{\angle}1|$ and $|d\text{W}x{\angle}2|$ are unordered with each other, and $|e\text{R}x{:}\{1,2\}|$ observes the values of both writes, hence causally depends on both. $|e\text{W}y{\angle}3|$ follows $e$’s read and hence causally depends on both writes as well. $|d\text{R}y{:}3|$ observes the result of $e$’s write and hence continues this causal dependency chain. In contrast, $|c\text{W}y{\angle}1|$ is dangling and has not been observed by any reader.
Interpreting A Partial Ordering. Assuming that we are designing a replication protocol atop a set of replica nodes, an intuitive way to interpret a partial ordering in the SOP model is to think from each replica’s perspective. Replicas are free to apply an arbitrary order between operations that are unordered with each other in the ordering decided by the consistency level. Figure 4 demonstrates this perspective.
@@
With a consistency level that always gives an SO ordering, all replicas agree on the same sequence of operations. With a consistency level that allows CPO or NPO ordering, replicas may apply operations in different orders, as long as everyone is coherent with the required validity constraints. This removes the need to coordinate a global sequence for operations that do not causally depend on each other, and is the root source of the scalability and availability benefits of causal+ and weaker consistency levels.
Why Causality. The causal property is desirable in many application scenarios. For example, (Lloyd et al., 2011) describes a scenario where client $c$ is sharing a photo with client $d$ by first uploading the photo to an image store $s$ and then adding a reference to the photo to the album $a$. Client $d$ then checks $c$’s album and, upon seeing a new reference, goes to fetch the referenced photo:
@@
Why Convergence. Compared to regular causal consistency, causal+ consistency demands a convergent conflict resolution mechanism for conflicting values observed. In other words, all read operations that observe the same set of unordered values on an object must resolve into the same return value. Examples of such conflict resolution mechanisms include last-writer-wins, taking-the-max, and taking-the-sum.
Without the convergence guarantee, regular causal consistency is allowed to forever return different values for reads on the same object from different clients. This may be undesirable in many application scenarios. For example, (Lloyd et al., 2011) describes a scenario where two clients, $c$ and $d$, happen to concurrently update the time for a reminder event $t$:
@@
Protocols & Systems. Causal dependency originates from causal memory models (Ahamad et al., 1995; Terry et al., 1994). It has been adopted by replication systems designed to address availability (Ladin et al., 1992; Bailis et al., 2013b; Bravo et al., 2021; Jiang et al., 2022; Belaramani et al., 2006; Petersen et al., 1997) and/or scalability (Lloyd et al., 2011; Almeida et al., 2013; Mehdi et al., 2017; Ferreira Loff et al., 2023; Belaramani et al., 2006; Petersen et al., 1997) concerns while preserving causality semantics.
@@ 4.4. Eventual Consistency
Eventual consistency, as the name suggests, is a consistency level that only requires reads on an object to return a consistent value if no updates are being made to the object (Vogels, 2008). There is no relationship constraint between operations, meaning that any pairs of operations issued by the same client are allowed to get reordered, let alone preserving causality, in the final ordering. Eventual consistency is widely adopted in geo-scale systems where the demand for high performance, scalability, and availability outweighs the need for timely consistency.
Eventual Convergence. Although eventual consistency is sometimes used interchangeably with weak consistency, it does impose one requirement on the service: the decided ordering must be convergent. In other words, after all the writers on an object become inactive and after all the writes become visible to readers, reads on the object must all return the same value. In our model, this is captured by the CPO constraint.
@@
Quiescent Consistency. A closely related, vaguely defined term is quiescent consistency, as mentioned in (Herlihy and Shavit, 2011). In a commonly accepted definition of quiescent consistency, special periods of physical time are identified, during which no write operations are happening. Every such contiguous time period is called a quiescence period. Each quiescence period orders all operations that are acknowledged ahead of the period before operations that start after the period. With this definition, quiescent consistency is weaker than eventual consistency, because it effectively makes no guarantees if a system-wide quiescence period never appears (Viotti and Vukolić, 2016).
Protocols & Systems. Eventual consistency is widely adopted by web-scale systems in the form of gossiping protocols and anti-entropy propagation (Schroeder et al., 1984; DeCandia et al., 2007; Elhemali et al., 2022; Lakshman and Malik, 2010).
@@ 4.5. Other Consistency Levels
In this section, we briefly describe the rest of the selected consistency levels other than the four most common ones. These levels explore different combinations of convergence and relationship constraints to refine the consistency level spectrum.
@@ 4.5.1. Regular Sequential Consistency
Helt et al. formalized the notion of regular sequential consistency in a recent work (Helt et al., 2021). Regular sequential consistency takes the middle ground between linearizability and sequential consistency. It combines the strengths of both by imposing different levels of relationship constraints for read-only operations and write operations. Specifically, all write operations (and RMWs) must honor the real-time property (denoted RT-W), while read operations are allowed to travel back in time as long as they still honor causality (denoted CASL-R).
@@
Invariant-equivalence to Linearizability. It is shown in (Helt et al., 2021) that regular sequential consistency is invariant-equivalent to linearizability, meaning that: 1) it is local (see Section 4.2 for the definition of implementation locality) and 2) it inherently supports RMW operations thanks to stable ordering of writes. However, it does not guarantee to capture external causality dependencies, making it still slightly weaker than linearizability. If external causality is not an issue, a linearizable replicated storage system can seamlessly adopt regular sequential consistency to improve the performance of read-only operations.
The transactional version of this consistency level is regular sequential serializability (Helt et al., 2021), where read-only transactions are allowed to get reordered in the serialization sequence, while all other transactions must honor RT. Similar properties have been exploited in transactional database systems that use certain Timestamp Ordering (T/O) optimistic concurrency control mechanisms (Yu et al., 2016).
@@ 4.5.2. Real-time Causal Consistency
The discussion around causality and convergence is tightly coupled with availability guarantees. Assuming that the system is composed of a set of symmetric message-passing replica nodes, Mahajan et al. have proven in (Mahajan et al., 2012) that real-time causal consistency is the strongest possible level that is achievable in an always-available, one-way convergent system (which is implied by our definition of sticky available in Section 5).
Fork-based Consistency Models. A family of fork-based consistency models has been developed to deal with Byzantine faults in a system containing untrusted replica nodes. For example, a fork-linearizable system ensures that if any two replicas have observed different orderings (i.e., forked by an adversary, even for one operation), then their writes will never be visible to each other afterward (i.e., they cannot be joined again). Fork causal consistency is a family of consistency levels that weaken causal consistency to tolerate Byzantine replicas and enforce causal consistency among correct replicas (Mahajan et al., 2011). It is proven that fork causal consistency is unachievable in an always-available, one-way convergent system, while bounded fork-join causal is achievable (Mahajan et al., 2012).
@@ 4.5.3. Causal Consistency
Section 4.3 has explained the most essential pieces of causal and causal+ consistency. To recap, a service that provides causal consistency must give an ordering that is NPO and CASL given any physical timeline workload. Such an ordering captures all the potential causality dependencies between operations. Compared to causal+, original causal consistency does not demand convergent conflict resolution, meaning that different clients are allowed to forever retrieve different values from reads on the same object.
@@ 4.5.5. PRAM Consistency
Pipeline Random Access Memory (PRAM) consistency (Lipton and Sandberg, 1988), or simply FIFO consistency, is a weaker consistency level than causal consistency, where causality across clients is not captured. It was originally defined within a shared memory system context. In our model of distributed replication, it is a consistency level that requires NPO and FIFO ordering.
@@ 4.5.6. Per-key Sequential Consistency
This consistency level was first studied in the PNUTS system (Cooper et al., 2008), a highly-concurrent data serving system that provides per-record consistency. However, modern distributed systems typically have complicated client-side logic layered on top of a non-transactional object store, where each client is interested in way more than one object. This makes the object-key-oriented consistency level less appealing than session-oriented causality levels. The photo-album case described in Section 4.3 would be a good example that demonstrates the limitations of per-key sequential consistency.

### v2

@@ 4.2. Sequential Consistency
External Causality Dependencies. So far we have assumed that all clients communicate only with the service and there are no external communication channels between clients that bypass the service, as depicted in Figure 1. However, in real distributed systems such as cloud databases (Kossmann et al., 2010; Verbitski et al., 2017; Dageville et al., 2016; Guo et al., 2022), clients of a replicated storage service may be part of a higher-level system. It is not uncommon for clients to coordinate with each other through external causality dependencies, which are impossible for the service to capture without preserving real-time dependencies.
In the example depicted by Figure 3, it could be that client $c$ first issues a write of value 1 to object $x$ and waits for its acknowledgment. It then sends a message to client $d$ through an external inter-client channel saying “I have finished my write to $x$ and you can go ahead to operate on $x$.” Client $d$ then issues its own write of value 2 and expects to read out 2 afterwards. However, since the message from $c$ to $d$ is external to the service, a sequentially consistent service may reorder $d$’s write ahead of $c$’s, and return value 1 for $d$’s read.
@@
A service that provides linearizability will be able to capture such implicit external dependencies because of the real-time property, as $|d\text{W}x{\angle}2|$ starts after $|c\text{W}x{\angle}1|$’s acknowledgment in physical time [^6] Note that this is not to be confused with the external consistency property in distributed transaction processing systems (Gifford, 1981; Corbett et al., 2013), which means that transactions are serialized into the same order as their commit order..
Implementation Locality. Herlihy and Wing have proven in (Herlihy and Wing, 1990) that a protocol that implements sequential consistency for each object individually does not necessarily guarantee overall sequential consistency across all operations. Formally, we say that sequential consistency is non-local: it is possible for an ordering to be SO and CASL on each object, while not SO or CASL overall.
@@
Adding support for read-modify-write (RMW) operations to such protocols is a non-trivial task (Burke et al., 2020). In particular, we cannot simply treat RMW operations in the same way as pure writes, because RMWs require a stable base value to determine the result of the read. Systems that demand compare-and-swap (CAS) operations (such as the LogOnce operation on shared logs (Guo et al., 2022)) may have to opt for a service that provides linearizability (or regular sequential consistency (Helt et al., 2021) as discussed in Section 4.5.1).
Protocols & Systems. Sequential consistency originates from memory consistency theory (Lamport, 1979; Adve and Gharachorloo, 1996; Hill, 1998). In the context of replicated objects, sequential consistency (or its per-key variant (Cooper et al., 2008)) is often seen in primary-backup systems (Hunt et al., 2010) and message streaming systems (Kreps et al., 2011; Redpanda, 2024; ZeroMQ, 2024) where writes may propagate to readable endpoints after acknowledgment. The transactional form, i.e., serializability (Berenson et al., 1995) plays an indispensable role in database systems.
@@ 4.3. Causal+ Consistency
If a global total order is not required, it may be desirable to further relax sequential consistency and embrace the family of causal consistency levels. Causal consistency stems from the definition of causal memory (Ahamad et al., 1995). Lloyd et al. pointed out in (Lloyd et al., 2011) that distributed replication protocols typically implement a slightly stronger version of causal consistency, which they term causal+ consistency. It is essentially causal consistency with convergent reads.
In our model, a service that provides causal+ consistency always gives an ordering that is CPO and CASL. Compared to sequential consistency, the ordering does not have to be a serial total order, but instead may leave certain operations from different clients unordered with each other. This opens up opportunities to improve the scalability of a replication protocol. However, all causal dependencies still have to be reflected in the decided ordering.
@@
Notice that $|c\text{W}x{\angle}1|$ and $|d\text{W}x{\angle}2|$ are unordered with each other, and $|e\text{R}x{:}\{1,2\}|$ observes the values of both writes, hence causally depends on both. $|e\text{W}y{\angle}3|$ follows $e$’s read and hence causally depends on both writes as well. $|d\text{R}y{:}3|$ observes the result of $e$’s write and hence continues this causal dependency chain, while $|c\text{W}y{\angle}1|$ is dangling and has not been observed by any reader.
Interpreting A Partial Ordering. Assuming that we are designing a replication protocol atop a set of replica nodes, an intuitive way to interpret a partial ordering in the SOP model is to think from each replica’s perspective. Replicas may each maintain a local ordering; different replicas are free to apply different orders for operations that are unordered from the global perspective. Figure 4 demonstrates this perspective.
@@
With a consistency level that always gives an SO ordering, all replicas agree on the same sequence of operations. With a consistency level that allows CPO or NPO ordering, replicas may apply operations in different orders, as long as everyone is coherent with the required validity constraints. This removes the need to coordinate a global sequence for writes that do not causally depend on each other, and is the root source of the scalability and availability benefits of causal+ and weaker consistency levels.
Why Causality. The causal property is desirable in many application scenarios. For example, COPS (Lloyd et al., 2011) describes a scenario where client $c$ is sharing a photo with client $d$ by first uploading the photo to an image store $s$ and then adding a reference to the photo to the album $a$. Client $d$ then checks $c$’s album and, upon seeing a new reference, goes to fetch the referenced photo:
@@
Why Convergence. Compared to plain causal consistency, causal+ consistency demands a convergent conflict resolution mechanism for conflicting values observed by a read. In other words, all read operations that observe the same set of unordered values on an object must resolve into the same return value. Examples of such conflict resolution mechanisms include last-writer-wins, taking-the-max, and taking-the-sum.
Without the convergence guarantee, causal consistency is allowed to forever return different values for reads on the same object from different clients. This is undesirable in many applications. For example, consider a scenario where two clients, $c$ and $d$, happen to concurrently update the time for a reminder event $t$ (Lloyd et al., 2011):
@@
Protocols & Systems. Causal dependency originates from causal memory models (Ahamad et al., 1995; Terry et al., 1994). It has been adopted by replication systems designed to address availability (Ladin et al., 1992; Bailis et al., 2013b; Bravo et al., 2021; Jiang et al., 2022; Belaramani et al., 2006; Petersen et al., 1997) and/or scalability (Lloyd et al., 2011; Almeida et al., 2013; Mehdi et al., 2017; Ferreira Loff et al., 2023; Belaramani et al., 2006; Petersen et al., 1997) concerns in large-scale cloud systems, while preserving useful causality semantics.
@@ 4.4. Eventual Consistency
Eventual consistency, as the name suggests, is a consistency level that only requires reads on an object to return a consistent value if no updates are being made to the object (Vogels, 2008). There is no relationship constraint between operations, meaning that any pair of operations issued by the same client are allowed to get reordered, let alone preserving causality, in the final ordering. Eventual consistency is widely adopted in geo-scale systems where the demand for high performance, scalability, and availability outweighs the need for timely consistency.
Eventual Convergence. Although eventual consistency is sometimes used interchangeably with weak consistency, it does impose one extra requirement on the service: the decided ordering must be convergent. In other words, after all the writers on an object become inactive and after all the writes become visible to readers, reads on the object must all return the same value. In our model, this is captured by the CPO constraint.
@@
Quiescent Consistency. A closely related, vaguely defined term is quiescent consistency (Herlihy and Shavit, 2011). In a commonly accepted definition, special periods of physical time are identified, during which no write operations are happening. Every such contiguous time period is called a quiescence period; all operations acknowledged ahead of the period are ordered before operations that start after the period. With this definition, quiescent consistency is weaker than eventual consistency, because it effectively makes no guarantees at all if a system-wide quiescence period never appears (Viotti and Vukolić, 2016).
Protocols & Systems. Eventual consistency is widely adopted by web-scale systems in the form of gossiping protocols and anti-entropy propagation (Schroeder et al., 1984; DeCandia et al., 2007; Elhemali et al., 2022; Lakshman and Malik, 2010). These systems value performance and scalability greatly and can tolerate inconsistencies.
@@ 4.5. Other Consistency Levels
In this section, we briefly describe the rest of the selected consistency levels other than the four most common ones. These levels explore different combinations of convergence and (variations of) relationship constraints to refine the consistency level hierarchy.
@@ 4.5.1. Regular Sequential Consistency
Helt et al. formalized the notion of regular sequential consistency in a recent work (Helt et al., 2021). It takes the middle ground between linearizability and sequential consistency. It combines the strengths of both by imposing different levels of relationship constraints for read-only operations versus write operations. Specifically, all writes (and RMWs) must honor the real-time property (denoted RT-W), while read operations are allowed to travel back in time as long as they still honor causality (denoted CASL-R).
@@
Invariant-equivalence to Linearizability. It is shown that regular sequential consistency is invariant-equivalent to linearizability (Helt et al., 2021), meaning that: 1) it is local (see Section 4.2) and 2) it inherently supports RMW operations thanks to stable ordering of writes. However, it does not guarantee to capture external causality dependencies, making it still slightly weaker than linearizability. If external causality is not an issue, a linearizable replication system can seamlessly adopt regular sequential consistency to improve the performance of read-only operations.
The transactional version of this consistency level is regular sequential serializability (Helt et al., 2021), where read-only transactions are allowed to get reordered in the serialized sequence, while all other transactions must honor RT. Similar properties have been exploited in transactional database systems that use Timestamp Ordering (T/O) optimistic concurrency control mechanisms (Yu et al., 2016).
@@ 4.5.2. Real-time Causal Consistency
Assuming that the system is composed of a set of symmetric message-passing replica nodes, Mahajan et al. have proven in (Mahajan et al., 2012) that real-time causal consistency is the strongest possible level that is achievable in an always-available, one-way convergent system (which is implied by our definition of sticky available in Section 5).
Fork-based Consistency Models. A family of fork-based consistency models has been developed to deal with Byzantine faults in a system containing untrusted replica nodes. For example, a fork-linearizable system ensures that if any two replicas have observed different orderings (i.e., forked by an adversary, even for one operation), then their writes will never be visible to each other afterwards (i.e., they cannot be joined again). Fork causal consistency is a family of consistency levels that weaken causal consistency to tolerate Byzantine replicas and enforce causal consistency among correct replicas (Mahajan et al., 2011).
@@ 4.5.3. Causal Consistency
Section 4.3 has explained causal and causal+ consistency. To recap, a service that provides causal consistency must give an ordering that is NPO and CASL given any physical timeline workload. Such an ordering captures all the potential causality dependencies between operations, but does not demand convergent conflict resolution, meaning that different clients are allowed to forever retrieve different values from reads on the same object.
@@ 4.5.5. PRAM Consistency
Pipeline Random Access Memory (PRAM) consistency (Lipton and Sandberg, 1988), or simply FIFO consistency, is a weaker consistency level than causal consistency, where causality across clients is not captured. It was originally defined for shared memory systems. In our framework, it is a consistency level that requires NPO and FIFO ordering.
@@ 4.5.6. Per-key Sequential Consistency
This consistency level was first studied in the PNUTS system (Cooper et al., 2008), a highly-concurrent data serving system that provides per-record consistency. However, modern distributed systems typically have complicated client-side logic layered on top of a non-transactional object store, where each client is interested in more than one object. This makes the object-key-oriented consistency level less appealing than session-oriented causality levels. The photo-album case described in Section 4.3 would be a good example that demonstrates the limitations of per-key sequential consistency.